OpenStudy (anonymous):

help?please... x^2-9x=0

jimthompson5910 (jim_thompson5910):

we can factor out the GCF x to get x^2 - 9x = 0 x*(x - 9) = 0

jimthompson5910 (jim_thompson5910):

from there, we use the zero product property to go from x*(x - 9) = 0 to x = 0 or x-9 = 0 the last step is to solve each equation for x. The first one is already done, so just solve x-9=0 for x to get x = 9 That means the two solutions are x = 0 or x = 9

jimthompson5910 (jim_thompson5910):

It's always a good idea to check your solutions Check x = 0 x^2 - 9x = 0 0^2 - 9*0 = 0 ... replace every x with 0 0 - 0 = 0 0 = 0 ... works ------------------ checking x = 9 x^2 - 9x = 0 9^2 - 9*9 = 0 ... replace every x with 9 81 - 81 = 0 0 = 0 ... works ------------------ So both solutions have been confirmed

OpenStudy (anonymous):

it is actually \[x ^{3}-9x=0\]

jimthompson5910 (jim_thompson5910):

so you'd have x(x^2 - 9) = 0 when you factor out the GCF x

jimthompson5910 (jim_thompson5910):

what do you get when you factor x^2 - 9 ?

OpenStudy (anonymous):

(x+3)(x-3)?

OpenStudy (anonymous):

nevermind i got it solved it like you did up there^ thanks!

jimthompson5910 (jim_thompson5910):

correct and you're welcome